AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:17 2025
Author
*
:
Subject
*
:
<p><strong>21:32</strong></p> <p>New run 59</p> <p>Right arm at 64.60 mm</p> <p>Left arm IN at 36.54 mm</p> <p>MBS working - hopefully</p> <p>Target 4.7E11</p> <p>Injection ca. 4E7</p> <p>horizontal beam position 8.75 mm</p> <p>vertical beam position -5mm</p> <p>Changed electron cooler set point to 244.85 V</p> <p>Read value 244.3 V</p> <p>Collector current 4.31A</p> <p>Histograms and statistics zero'd. Code reset</p> <p>Rates OK. Temps OK.</p> <p><strong>22:44</strong></p> <p>Screenshots of the above</p> <p> </p> <p><strong>23:00</strong></p> <p>Detector 0 shows usual strong noise. Shook twice. Fixed it. Beam loss in correspondence with this event.</p> <p>Beam mostly recovered a few minutes later.</p> <p><strong>23:28</strong></p> <p>Detector 0 shows noise. Shook twice with servo. No result. Shook with pneumatic. Fixed. Some beam instabilities again.</p> <p><strong>01:00</strong></p> <p>Temps OK. Rates OK.</p> <p>Merger OK. Write to disk OK (file 59_13)</p> <p><strong>01:36</strong></p> <p>Detector 0 shows usual strong noise. Shook it. Fixed it.</p> <p><strong>01:46</strong></p> <p>Detector 0 shows noise. Shook twice with servo. No result. Shook with pneumatic. Fixed.</p> <p>The noise increase is definitely correlated with spikes in YRT1DT1MK current reading. Unclear what the reading is.</p> <p><strong>02:36</strong></p> <p>Target density has been consistentyl above 1E12 recently, very unstable and trending forward. Increased Lakeshore setpoint by 0.5 K.</p> <p>Density decreases to high 1E11, and appear significantly more stable. E1 pressure reading also decreasing to 1E-1 mbar.</p> <p>Temps OK. Stats OK. Merger OK. Write to disk OK.</p> <p><strong>03:58</strong></p> <p>Temps OK. Stats OK. Merger OK. Write to disk OK.</p> <p>Target still unstable and now a bit too low density. Trying to slowly lower the temperature in 0.1 K steps.</p> <p><strong>04:55</strong></p> <p>Temps OK. Stats OK. Merger OK. Write to disk OK.</p> <p>Target around 6E11 atoms/cm2, with Lakeshore temperature at 30.8 K. Waiting for it to stabilise.</p> <p><strong>05:34</strong></p> <p>Detector 0 noisy. Shook. Fixed it.</p> <p>Realised I had not moved the detector back at 01:46 after shaking it. Most of this run has been taken with detector 0 fully out, i.e. 18.86 mm.</p> <p>Moving in and moving to a new run.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown